Lupita has five times as much money as David. Lupita also has $20 more than David. How much does each person have?

Answer by checkley79(3341)

L=5D
L=D+20
5D=D+20
5D-D=20
4D=20
D=20/4
D=$5.00 AMOUNT DAVID HAS.
L=5*5=$25.00 AMOUNT LIPITA HAS.
PROOF:
25=5+20
25=25